# Some remarks on real algebraic maps which are topologically special generic maps and generalize the canonical projections of the unit spheres

### v5

@@ Abstract.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the corresponding number in the open interval $(0,1)$.
@@
Thus we have a homeomorphism from the smooth manifold $M_{0}$ onto a smooth manifold $M$. The restriction of this to the set $M_{0}-(({\partial\bar{N_{0}}})\times\{0\})$ gives a diffeomorphism.
We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m-n+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$. The resulting map is also special generic by the construction. It is also product-organized. We define a special generic map $f$ as this map. An essentially same argument also appears in Main Theorems 1 and 2 and their proof.
@@
We use the function $g$ before. We can map each point of the form $(x,y)\in({\rm Int}\ {\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1})\bigcap W_{0}$ to $(x,g(t){\Phi}_{F_{S^{m-n}},a}({\rm Pr}_{{\Sigma}^{m-n},a,1}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y),\frac{{\rm Pr}_{{\Sigma}^{m-n},a,2}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y)}{(\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T})}\times T_{0}))$ according to the following rule in the case $y$ is not the origin.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the corresponding number in the interval $(0,1)$.
@@
This rule is in fact same as the previous one. This gives a diffeomorphism. We map each point of the form $(x,0)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. Thus we have a homeomorphism from $W_{0}$. The image of the homeomorphism is the disjoint union of the image of the previous diffeomorphism and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$.
From this we can see that the set $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$. The projection is given by the restriction of the canonical projection ${\pi}_{m+1,n}$. In other words, it is also regarded as a ”canonical projection of $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ onto ${\rm Int}\ {\bar{N}}_{0}$”.
@@
for $0<t^{\prime}\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$. We present several properties based on our presented arguments. We can also know $M_{0,t^{\prime}}:=\{(x,y)\in\bar{N_{0}}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t^{\prime})=0\}$. For distinct real numbers $t^{\prime}=t_{1},t_{2}$ here, the sets $M_{0,t_{1}}-((\partial\bar{N_{0}})\times\{0\})$ and $M_{0,t_{2}}-((\partial\bar{N_{0}})\times\{0\})$ are disjoint. These sets $M_{0,t^{\prime}}-((\partial\bar{N_{0}})\times\{0\})$ are $m$-dimensional smooth manifolds with no boundaries. They are seen as the graphs of smooth functions as $M_{0}-((\partial\bar{N_{0}})\times\{0\})$ and diffeomorphic to $M_{0}-((\partial\bar{N_{0}})\times\{0\})$. Here some variable $y_{j}$ is regarded as the variable of the target of the function. We can show these facts by arguments similar to the corresponding arguments in the case of $M_{0}$ and $M_{0}-((\partial\bar{N_{0}})\times\{0\})$. We can also know that the manifold $M_{0,t^{\prime}}$ is an $m$-dimensional smooth and closed manifold and a non-singular real algebraic manifold. This manifold is also homeomorphic to the manifold $M_{0}:=M_{0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}}$.
@@
From our definitions, we also have the relations $W_{0}=({\bar{N}}_{0}\times\{0\})\bigcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}M_{0,t^{\prime}})$ and $W_{0}=({\bar{N}}_{0}\times\{0\})\sqcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}(M_{0,t^{\prime}})-((\partial{\bar{N}}_{0})\times\{0\}))$.
@@
We go back to the trivial bundle $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ over ${\rm Int}\ {\bar{N}}_{0}$ which is not seen as a smooth bundle. This is also regarded as a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). For this, first, remember that the projection is the restriction of the canonical projection ${\pi}_{m+1,n}$. Remember that around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function by regarding some variable $y_{j}$ as the variable of the target of the function. Remember also that the manifold $M_{0,t^{\prime}}$ enjoys a similar property, for example. The relation $S(f_{0})=(\partial{\bar{N}}_{0})\times\{0\}$ also holds.
@@
In the case $F_{S^{m-n}}(y,t)=||y||^{2}-t$, the previous smooth bundle $W_{0}-S(f_{0})$ which may not be trivial in general is also a trivial smooth bundle. Our arguments here generalize ones from STEP 3 in the proof of Main Theorem 2.
Last we consider a small isotopy as in the proof of Main Theorems 1 and 2. This completes the proof of the statement of Main Theorem 1.
@@
We can also check the property (5) by investigating arguments of STEP 3 of our proof of Main Theorems 1 and 2 and ”Discussion 14” of [21] again. In other words, we can check that around the set $S(f_{0})$, $f_{0}$ is represented as the product map of a natural function and the identuty map on $S(f_{0})$ as the case of ”$f_{0}$ in Main Theorems 1 and 2”.
@@
We can also check Main Theorem 3 (6) easily thanks to our arguments. A result close to this is also presented in Theorem 1 (3) of [18]. There ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$. The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there. Only exposition on ”homotopy” is presented there.
@@
Note that $F_{S^{m-n}}(y,t)=t-||y||^{2}$ gives one of simplest cases where $y\in{\mathbb{R}}^{m-n+1}$. For this we can choose $a$ as an arbitrary positive real number. The bundle $M_{{\rm B},0}$ in (1) is regarded as a trivial linear bundle: ${\Delta}^{k}$ is $D^{k}$ in this case. Furthermore, in this case, we can consider a suitable diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}$ and regard that (2) can be discussed in the smooth category, for example. In short, we can argue all in the smooth situation in Main Theorem 3. This is also for Main Theorem 2. Main Theorem 2 gives one of most important specific cases.
This is generalized for more general cases with $m-n\neq 4$ for example. This is thanks to fundamental arguments on singularity theory and differential topology. This is also thanks to well-known classical and sophisticated theory of diffeomorphisms on spheres. This is also related to studies on diffeomorphism groups of unit spheres, studied in [12], and so-called ”pseudo-isotopy”, pioneered in [1], for example. We also apply related arguments implicitly in some of our previous arguments in the present paper.
See also Remark 4, in the next section, for example.
We give additional explicit cases as Examples. They are also a kind of remarks.
@@
  • The function $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ is a Morse function.
  • The restriction of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ to its zero set is a Morse function on a smooth closed manifold. The manifold is a non-singular real algebraic manifold of dimension $m-n$.
@@
We also note that Main Theorem 3 respects the case $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y):={\Sigma}_{j=1}^{m-n+1}k_{j,1}{y_{j}}^{2k_{j,2}}$ with positive numbers $k_{j,1}$ and positive integers $k_{j,2}$, for example. See also Remark 3 of [15]. Some examples of Example 2 are regarded as specific cases of the present cases.
3. Meanings of our main result and remarks.
The first remark compares our main result to previous results by the author.
@@
In [15, 16, 18], we consider cases where polynomials are given beforehand. In short, only STEP 2 in our proof of Main Theorems 1 and 2 is considered and related short exposition on STEP 3 in our proof of Main Theorems 1 and 2 is presented where the latter presentation is a kind of short remark. This does not affect main ingredients of [15, 16, 18]. STEP 3 explicitly shows that $f_{0}$ is ”topologically special generic” first. These studies also consider explicit cases. For example, polynomials are of degree $2$ and for spheres. These studies are regarded as studies finding nice classes of real algebraic maps including the canonical projections of unit spheres as simplest cases.
@@
As related studies, we also present a comment on another problem, studying global structures of functions obtained by composing canonical projections. Some are closely related to Remark 1.
More precisely, such studies are on Reeb graphs of these functions. The Reeb graph of a smooth function is the graph obtained as the quotient space of the manifold of the domain and consists of connected components of all preimages. The vertex set of the Reeb graph consists of all elements representing connected components containing some singular points of the function.
[41] is a pioneering study considering Reeb graphs. Such tools are fundamental and strong tools in studying manifolds by using graphs. In 2006, [45] has asked whether we can reconstruct a (nice) smooth function on a suitable manifold whose Reeb graph is isomorphic to a prescribed graph. [24, 34] show explicit answers. [14] is motivated by such studies. There the author has succeeded in construction of nice smooth functions with prescribed preimages first.
[15] considers this problem in the real algebraic category and constructs real algebraic functions with prescribed Reeb graphs in explicitly and nicely considered situations first. This is also regarded as a study finding a class of real algebraic functions including the canonical projection of a unit sphere. For the canonical projection of the unit sphere, its Reeb graph consists of exactly one edge and exactly two vertices. For this, [4, 49, 50] are also closely related, help the author to do such studies, and motivate us to study further. They present some important topics on singularity theory, real algebraic geometry and topology of non-singular real algebraic curves in ${\mathbb{R}}^{2}$ and regions in ${\mathbb{R}}^{2}$ surrounded by non-singular real algebraic curves.
Remark 3 is a remark on STEP 1 in our proof of Main Theorems 1 and 2. This is also independent of our present study. Here we also add a remark on singularity theory of smooth maps, Mather has pioneered and studied mainly in [26, 27, 28, 29, 30, 31, 32, 33]. See also [42]. Remark 3 is also closely related to [4]. This is on graphs in ${\mathbb{R}}^{2}$ which are generic from the viewpoint of singularity theory. This mainly studies regions surrounded by non-singular real algebraic curves and collapsing naturally to the graphs. We do not present elementary notions on graphs precisely. We except related knowledge. We also note that we omit several rigorous discussions in Remark 3 and that for our main ingredients of the present paper we do not need to understand this rigorously. However, we give comments that this is also related to Remark 1 and 2.
@@
An algebraic domain is an open set in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint non-singular connected real algebraic curves. The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of all connected components of all preimages for the function ${\pi}_{2,1}{\mid}_{\overline{R}}$: the natural equivalence relation is considered. Here we also consider a situation such that the vertex set of the graph can be defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\partial\overline{R}}$. The paper says that for the embedded graph $e(K)$, we can have an algebraic domain $R_{K}$ enjoying the following.
@@
      $\partial\overline{R_{K}}$ is regarded as the intersection of some bounded open set and the zero set of a real polynomial.
A desired region $R_{K}$ can be also obtained by an approximation of the Weierstrass type in the class $C^{2}$. According to arguments similar to ones in ”Special Case 5” of [21], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial.
The following is a remark on Main Theorem 3.
@@
We can also show Main Theorem 3 in the case $m=n$ by respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively. In this case, $M$ and $M_{0}$ are diffeomorphic to the so-called double of $\overline{N_{0}}$.
@@
Construction of real algebraic curves in the real affine space ${\mathbb{R}}^{2}$ and real algebraic hypersurfaces has been difficult and important in real algebraic geometry. This comes from Hilbert’s 16th problem and for example, Viro’s sophisticated theory such as [10, 54] presents strong tools. At present, even explicit construction and classifications for cases of real algebraic curves are difficult. Our pioneering studies of the construction respect unit spheres and canonical projections of them where we do not restrict attention to $1$-dimensional cases. Our pioneering studies are also new in concentrating not only on the hypersurfaces themselves but also on their canonical projections.
@@
We close our paper by a short remark on [18].
Remark 6.
[18] has no essential errors and we have also found that we need several more rigorous exposition. Some of these important arguments are discussed in the present paper. We will revise the version and submit our revised version to the refereed journal, based on the presented comments by editors and referees on [18] and our new arguments and idea in the present paper. Note also that the content of [18] is independent of the content of the present paper essentially. We have also found we can revise [18] in a self-contained way although we need to respect some of our paper. We will do so as we can.

### v6

@@ Abstract.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the value of the projection to the 2nd component of $x$ here.
@@
Thus we have a continuous bijection ${\Phi}^{\prime}$ between compact and connected manifolds, which is also a homeomorphism ${\Phi}^{\prime}$ from the smooth manifold $M_{0}$ onto another smooth manifold $M$.
We can also have the composition of the canonical embedding $M\subset{\mathbb{R}}^{m+1}$ with the canonical projection ${\pi}_{m+1,n}:{\mathbb{R}}^{m+1}\rightarrow{\mathbb{R}}^{n}$. The resulting map is also special generic by the construction. It is also product-organized. We define a special generic map $f^{\prime}:M\rightarrow{\mathbb{R}}^{n}$ as this map. Note that this is not real algebraic (in general).
An essentially same argument also appears in Main Theorems 1 and 2 and their proof. We further present another argument, an argument similar to which also appears there.
@@
We use the function $g$ before. We can map each point of the form $(x,y)\in({\rm Int}\ ({\bar{N}}_{0})\times({\mathbb{R}}^{m-n+1}-\{0\}))\bigcap W_{0}$ to $(x,g(t){\phi}_{\Sigma}\circ{\rm Pr}_{{\Sigma}^{m-n},a}\circ{{\Phi}_{F_{S^{m-n}},a}}^{-1}(y))$ according to the following rule.
  • If $x\in N(\partial{\bar{N}}_{0})-\partial N(\partial{\bar{N}}_{0})=(\partial{\bar{N}}_{0})\times(0,1)$, then $t$ is the value of the projection to the 2nd component of $x$ here.
@@
This rule is in fact same as the previous one. This gives a continuous bijection. We map each point of the form $(x,0)\in{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$ to $(x,0)$. Thus we have a continuous bijection from $W_{0}$ to another compact and connected manifold, extending the homeomorphism ${\Phi}^{\prime}$ to a new homeomorphism. The image of the homeomorphism is the disjoint union of the image of the homeomorphism just before and ${\bar{N}}_{0}\times\{0\}\subset{\bar{N}}_{0}\times{\mathbb{R}}^{m-n+1}$. For here, remember again the condition (3), assuming ${F_{S^{m-n},{\mathbb{R}}^{m-n+1}}}^{-1}(0)=\{0\}\subset{\mathbb{R}}^{m-n+1}$.
From this, we can see that the set $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ is regarded as the total space of a trivial bundle over ${\rm Int}\ {\bar{N}}_{0}$ which may not be a smooth bundle and whose fiber is homeomorphic to $D^{m-n+1}$. The projection is given by the restriction of the canonical projection ${\pi}_{m+1,n}$. In other words, it is also regarded as a ”canonical projection of $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ onto ${\rm Int}\ {\bar{N}}_{0}$”.
@@
for $0<t^{\prime}\leq\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}$. We present several properties based on our presented arguments. We can also know $M_{0,t^{\prime}}:=\{(x,y)\in\bar{N_{0}}\times{\mathbb{R}}^{m-n+1}={\mathbb{R}}^{m+1}\mid F_{S^{m-n}}(y,t^{\prime})=0\}$. For distinct real numbers $t^{\prime}=t_{1},t_{2}$ here, the sets $M_{0,t_{1}}-((\partial\bar{N_{0}})\times\{0\})$ and $M_{0,t_{2}}-((\partial\bar{N_{0}})\times\{0\})$ are disjoint. These sets $M_{0,t^{\prime}}-((\partial\bar{N_{0}})\times\{0\})$ are $m$-dimensional smooth manifolds with no boundaries. They are seen as the graphs of smooth functions as $M_{0}-((\partial\bar{N_{0}})\times\{0\})$ and diffeomorphic to $M_{0}-((\partial\bar{N_{0}})\times\{0\})$. Here some variable $y_{j}$ is regarded as the variable of the target of the function. We can also know that the manifold $M_{0,t^{\prime}}$ is an $m$-dimensional smooth and closed manifold and a non-singular real algebraic manifold. This manifold is also homeomorphic to the manifold $M_{0}:=M_{0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}}$. We can show these facts by arguments similar to the corresponding arguments in the case of $M_{0}$ and $M_{0}-((\partial\bar{N_{0}})\times\{0\})$.
@@
From our definitions, we also have the relations $W_{0}=({\bar{N}}_{0}\times\{0\})\bigcup({\bigcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}M_{0,t^{\prime}})$ and $W_{0}=({\bar{N}}_{0}\times\{0\})\sqcup({\sqcup}_{t^{\prime}\in(0,\frac{{\prod}_{j=1}^{l}(f_{{\partial}_{j}}(x))}{T}]}(M_{0,t^{\prime}}-((\partial{\bar{N}}_{0})\times\{0\})))$.
@@
We go back to the trivial bundle $W_{0}-((\partial{\bar{N}}_{0})\times\{0\})$ over ${\rm Int}\ {\bar{N}}_{0}$ which is not seen as a smooth bundle. This is also regarded as a smooth bundle by so-called Ehresmann’s fibration theorem (or a relative version of this). For this, remember that around a point of the form $(x,y)\in(({\rm Int}\ {\bar{N}}_{0})\times{\mathbb{R}}^{m-n+1})\bigcap M_{0}$, it is represented as the graph of a canonically obtained smooth function by regarding some variable $y_{j}$ as the variable of the target of the function. Remember also that the manifold $M_{0,t^{\prime}}$ enjoys a similar property, for example.
@@
In the case $F_{S^{m-n}}(y,t)=t-||y||^{2}$, the previous smooth bundle $W_{0}-S(f_{0})$ which may not be trivial in general is also a trivial smooth bundle. Our arguments here generalize ones from ”STEP 3 in the proof of Main Theorems 1 and 2”.
These arguments also imply the relation $f_{0}\circ{\Phi}^{\prime}=f^{\prime}$. We can also discuss for the existence of the homeomorphisms of (6) similarly. More precisely, to obtain a product-organized special generic map $f:M\rightarrow{\mathbb{R}}^{n}$ satisfying the relation $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}$ and a homeomorphism $\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ satisfying the relation $f^{\prime}=\phi\circ f$ and $\phi(f(S(f)))=f_{0}(S(f_{0}))$, we apply Proposition 3 with Remark 1 (2): remember also that the embeddings ${\bar{f}}_{{\mathbb{R}}^{n}}$ and ${\bar{f}}_{{\mathbb{R}}^{n},0}$ are mutually isotopic smooth embeddings. In addition, the homeomorohism ${\Phi}^{\prime}$ is our desired homeomorphism $\Phi$. This completes the proof of the statement similar to that of Main Theorem 1 with the properties (5, 6).
@@
We can also check the property (7) easily by our construction. We can check that around the set $S(f_{0})$, $f_{0}$ is represented as a natural product map as the case of ”$f_{0}$ in Main Theorems 1 and 2”.
@@
We can also check Main Theorem 3 (8) easily by our construction. A result close to this is also presented in the previous version (2nd version) of [18] ([18, Theorem 1 (3)]). There ”$\overline{D^{\prime}}$” is seen as the ($m+1$)-dimensional bounded and compact manifold $W_{0}$ bounded by $M_{0}$. The topology of the ($m+1$)-dimensional manifold is not so precisely discussed there. Only exposition on ”homotopy” is presented there. Of course we do not assume arguments and result of [18].
@@
The case $F_{S^{m-n}}(y,t)=t-||y||^{2}$ ($y\in{\mathbb{R}}^{m-n+1}$) gives one of simplest cases. For this we can choose $a$ as an arbitrary positive real number. The bundle $M_{{\rm B},0}$ in (1) is regarded as a trivial linear bundle with ${\Delta}^{k}=D^{k}$ in this case. Furthermore, in this case, we can consider a suitable diffeomorphism ${\Phi}_{F_{{S}^{m-n}},a}$ and regard that (2) can be discussed in the smooth category, for example. In short, we can argue all in the smooth situation in Main Theorem 3 except arguments related to the deformation of the embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ of STEP 1. This is also for Main Theorems 1 and 2. Main Theorems 1 and 2 give one of most important specific and simplest cases.
This is generalized for more general cases with $m-n\neq 4$ for example. This is thanks to fundamental arguments on singularity theory and differential topology. This is also thanks to well-known classical and sophisticated theory of diffeomorphisms on spheres. This is also related to studies on diffeomorphism groups of the unit spheres, studied in [12], and so-called ”pseudo-isotopy”, pioneered in [1], for example.
We also apply related arguments implicitly in some of our previous arguments in the present paper. See also Remark 5, in the next section, for example.
We give additional explicit cases as Examples. They are also a kind of our new remark.
@@
  • The restriction of $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}$ to its zero set is a Morse function on an ($m-n+1$)-dimensional smooth manifold with no boundary.
@@
  • The preimage $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(t)$ is a non-singular real algebraic manifold of dimension $m-n$ for $t>0$ and diffeomorphic to $S^{m-n}$.
@@
We also note that Main Theorem 3 respects the case $F_{S^{m-n},{\mathbb{R}}^{m-n+1}}(y):={\Sigma}_{j=1}^{m-n+1}k_{j,1}{y_{j}}^{2k_{j,2}}$ with positive numbers $k_{j,1}$ and positive integers $k_{j,2}$, for example. See also [15, Remark 3]. Some examples of Example 2 are regarded as specific cases of the present cases.
3. Additional comments on our new result.
Remark 2 compares our main result to previous results by the author.
@@
In [15, 16, 18], we consider cases where in Main Theorems, we do not need to change the embedding ${\bar{f}}_{{\mathbb{R}}^{n}}$ and can regard ${\bar{f}}_{{\mathbb{R}}^{n}}={\bar{f}}_{{\mathbb{R}}^{n},0}$ with the polynomials $f_{{\partial}_{j}}$ being given beforehand. In short, only ”STEP 2 in our proof of Main Theorems 1 and 2” is considered and related short exposition on ”STEP 3 in our proof of Main Theorems 1 and 2” is presented there. The latter presentation is a kind of short remark. This does not affect main ingredients of [15, 16, 18]. ”Our STEP 3” explicitly shows that $f_{0}$ is ”topologically special generic” first.
These studies [15, 16, 18] also consider very explicit cases. For example, the polynomials are of degree $2$ and for spheres. These studies are also regarded as studies finding nice classes of real algebraic maps including the canonical projections of the unit spheres as simplest cases.
@@
As related studies, we also present a comment on another problem, studying global structures of functions obtained by composing the canonical projections to the real algebraic maps in Main Theorems. Some are closely related to Remark 2, just before.
More precisely, such studies are on Reeb graphs of these functions. The Reeb graph of a smooth function is the graph obtained as the quotient space of the manifold of the domain and consists of all connected components of all preimages of single points. The vertex set of the Reeb graph consists of all elements representing connected components containing some singular points of the function.
[41] is a pioneering study considering Reeb graphs. Such tools are fundamental and strong tools in studying manifolds by using graphs. In 2006, [45] has asked whether we can reconstruct a (nice) smooth function on a suitable manifold whose Reeb graph is isomorphic to a prescribed graph. [24, 34] also show explicit answers. [14] is motivated by such studies. There the author has succeeded in construction of nice smooth functions with prescribed preimages, first.
[15] considers this problem in the real algebraic category and constructs real algebraic functions with prescribed Reeb graphs in explicitly and nicely considered situations, first. This is also regarded as a study finding nice classes of real algebraic functions including the canonical projection of the unit sphere into $\mathbb{R}$. For the Reeb graph of the canonical projection of the unit sphere of dimension at least $2$ into $\mathbb{R}$, its Reeb graph consists of exactly one edge and exactly two vertices. For this, [4, 49, 50] are also closely related, help the author to do such studies, and motivate us to study further. They also present some important topics on singularity theory, real algebraic geometry and topology of non-singular real algebraic curves in ${\mathbb{R}}^{2}$ with regions surrounded by these curves in ${\mathbb{R}}^{2}$.
Remark 4 is on ”STEP 1 in our proof of Main Theorems 1 and 2”. This is also independent of our present study. Here we also give comments on singularity theory of smooth maps, Mather has pioneered and studied mainly in [26, 27, 28, 29, 30, 31, 32, 33]. See also [42]. Remark 4 is also closely related to [4]. This is on graphs in ${\mathbb{R}}^{2}$ which are generic from the viewpoint of singularity theory. This mainly studies regions surrounded by non-singular real algebraic curves and collapsing naturally to the graphs. We do not present elementary notions on graphs precisely. We except related elementary knowledge. We also note that we omit several rigorous discussions in Remark 4 and that for our main ingredients of the present paper, we do not need to understand this rigorously. We also give comments that this is also related to Remarks 2 and 3.
@@
An algebraic domain is an open set in ${\mathbb{R}}^{2}$ surrounded by mutually disjoint non-singular connected real algebraic curves. The Poincaré-Reeb graph of an algebraic domain $R$ is the graph defined as the quotient space of the closure $\overline{R}$ and the space consisting of all connected components of all preimages of all single points for the function ${\pi}_{2,1}{\mid}_{\overline{R}}$: the natural equivalence relation on the closure $\overline{R}$ is defined here. The vertex set of the graph can be and is defined as the set of all points for connected components containing some singular points of the function ${\pi}_{2,1}{\mid}_{\partial\overline{R}}$. The paper says that for the embedded graph $e(K)$, we can have an algebraic domain $R_{K}$ enjoying the following.
@@
      $\partial\overline{R_{K}}$ is regarded as the intersection of some bounded open set and the zero set of some real polynomial function.
Our desired region $R_{K}$ can be also obtained by a kind of approximation of the Weierstrass type in the class $C^{2}$. According to arguments similar to ones in [21, Special Case 5], we can consider the $C^{2}$ approximation here in such a way that our $\partial\overline{R_{K}}$ here is regarded as the zero set of another real polynomial function.
The following is on Main Theorem 3 and related to some theory of diffeomorphism groups and differential topolgy of manifolds.
@@
In Main Theorem 3, let $m=n$. By respecting the well-known facts that the diffeomorphism group of $D^{1}$ has the homotopy type of the two-point discrete set and that each point of the two-point set corresponds to orientation preserving diffeomorphisms and orientation reversing ones, respectively, $M$ and $M_{0}$ are shown to be diffeomorphic to the so-called double of the manifold $\bar{N_{0}}$.
@@
Construction of real algebraic curves in the real affine space ${\mathbb{R}}^{2}$ and real algebraic hypersurfaces of a general dimension has been difficult and important in real algebraic geometry. This comes from Hilbert’s 16th problem and for example, Viro’s sophisticated theory such as [10, 54] presents strong tools. At present, even explicit construction and classifications for cases of real algebraic curves are difficult. Our pioneering studies of the construction respect the unit spheres and the canonical projections of them and we do not restrict attention to the $1$-dimensional case. Our pioneering studies are also new in concentrating not only on the hypersurfaces themselves but also on their canonical projections.